LHC 2010 - latest news

Week 43: Machine coordinators: Mike Lamont – Ralph Assmann

Main aims of the week:

LHC coordination vistar



Link to report from morning meeting

Sunday 31st October

  • 00h47: Start ramp. 109b. 50ns. ~1.1e11 p / bunch.
  • 02h24: Stable beams. Fill 1459. First collisions with 50ns bunch spacing.
  • 07h30: Adjust. End-of-fill studies.
  • 10h10: Beam dump by operation. Ramp down and recycle.
  • 11h12: Injection.
  • 12h12: 108b (1 batch 12b + 4 batches 24b, 50ns) dumped with interlock from vacuum. Warm-cold transition on Q6.L7. Valves closed. See info in linked file.
  • 12h15: Go for scrubbing with 84b (1 batch 12b + 3 batches 24b, 50ns).

Slides from 9:00 meeting


Vacuum interlock at 12h12



Saturday 30th October

  • Overnight Friday to Saturday: successful completion of TOTEM run following by Alice length scale calibration; longitudinal scan; abort gap filling and cleaning studies at 3.5 TeV.
  • During day somewhat perturbed second pass damper, RF and injection studies
  • 08h07: Abort gap filling study.
  • 09h48: Beam dump by operation.
  • 13h45: SPS longitudinal blow-up set up.
  • 13h52: QPS A12 communication problem
  • 14h08: QPS OK.
  • 14h24: ATLAS access.
  • 16h20: Injection pilots.
  • 16h45: Injection setup 50ns. IP6 interlock test. RF setup 50ns.
  • 21h26: ATLAS access.
  • 22h44: ATLAS access.
  • 23h49: Injection.

Slides from 9:00 meeting

Friday 29th October

  • 00h36: Start ramp. 368b.
  • 01h27: Beam dump at 3.5 TeV. BLM trigger in Q4.L2 (UFO?). Losses start in injection kicker.
  • 02h35: Injection.
  • 03h15: Start ramp. 368b.
  • 04h16: Stable beams. Fill #1453. Turn-around: 2h40.
  • 08h30: Length scale calibration in ALICE.
  • 11h00: Beam dump by operations.
  • 50 ns commissioning: transverse damper, injection
  • 21:00 Start TOTEM run

Slides from 8:30 meeting


Ufino as seen by Alice

c/o Jan Fiete Gross-Oetring


Altas halo bct vacuum

c/o Witold Kozanecki

Thursday 28th October

  • 00h37: Stable beams. Fill #1450.
  • 15h17: Beam dump by operations. Integrated: 6.4/6.2/5.2 pb-1 (ATLAS/CMA/LHCb). Length of fill: 14h40.
  • 15h25: Ramp down. Cycle. Change of monitor factor for 4 BLM’s (0.3 -> 0.5).
  • 16h45: Injection.
  • 19h46: Start ramp. 368b.
  • 20h12: Beam dump just at flat top. ALICE removed beam permit. ALICE BCM triggered.
  • 22h51: Injection.

Slides from 8:30 meeting

Wednesday 27th October

  • 01h04: Both beams filled with 424b. Had to dump at 450 GeV to reset QPS_OK on RB_A23/RQF/RQD_A23. Ramp down and pre-cycle.
  • 02h46: Injection.
  • 04h06: Start ramp. 424b.
  • 04h23: Beam dump just after arriving to 3.5 TeV. TCTH.4R8.B2 BLM in IR8 triggered on 0.6s running sum. A few 10’s seconds before saw strange loss patterns. Beam-gas scattering?
  • 05h26: Injection.
  • 07h21: Start ramp. 364b.
  • 07h27: Beam dump due to SIS error on beam energy (timing telegram or something else?). Beam dump fine.
  • 07h40: CMS start ramp-down.
  • 09h00: Start of MD on abort gap cleaning in shadow of CMS work.
  • 16h00: Injection. CMS back.
  • 18h03: Injection kicker flashover IR2.
  • 18h10: Water leak CMS – emergency access.
  • 19h35: Injection.
  • 22h46: Start ramp. 364b.

Slides from 8:30 meeting


Slides from LMC

Tuesday 26th October

  • 00h54: Start ramp, 368b.
  • 01h18: Beam dump due to UFO after 5 min at 3.5 TeV.
  • 01h26: Ramp down, recover.
  • 02h35: Injection.
  • 04h36: Start ramp, 368b.
  • 05h34: Stable beams. Fill #1443.
  • 07h50: Beam dump due to UFO Q8.L5.
  • 07h55: Ramp down and recover.
  • 08h42: Injection. SPS transmitter problem (9h39-10h08).
  • 12h26: Start ramp, 368b.
  • 13h35: Stable beams. Fill #1444.
  • 16h56: Vacuum solenoid off/on.
  • 20h48: Beams dumped by EIC.
  • 21h14: Increase of various BLM thresholds through monitor factor.
  • 23h18: Interlock beam dump test with 324b. OK.

Slides from 8:30 meeting

Monday 25th October

  • 02h35: Stable beams, 368b, fill #1440.
  • 13h55: Adjust mode. End of fill study on beam tails.
  • 14h12: Beam dump, ramp down and recover.
  • 15h41: Injection.
  • 18h42: Start ramp, 368b.
  • 19h03: Problem during setting of crossing angle. Beam dumped by EIC following procedure.
  • 19h03-20h07: Cryo communication problem S34.
  • 21h32: Injection.

Slides from 8:30 meeting

Serge Claude presentation